Grade 11General Physics

A 10 m long wire of uniform cross-section and 20 ohm resistance is used in a potentiometer. The wire is connected in series with a battery of 5 V along with an external resistance of 480 ohm. If an unknown emf E is balanced at 6.0 m length of the wire, calculat

tanisha singh
12 Years agoGrade 11

1 Answer

Sagar Chawla
9 Years ago
Net resistance in external circuit =(480+20)ohm=500ohmWe know:-I=E/R=5/500=0.01APotential across the wire=IR=0.01×20=0.2THEREFORE,Potential Gradient=0.2/10=0.02V/mE=Potential Gradient×length of the wire where it gets balanced=0.02×6=0.12V